A 1500-kg car is ascending the ramp of a garage inclined at 10° with constant speed when the force applied is P. determine the magnitude of force P if u=0.70
1
Expert's answer
2018-08-13T12:00:09-0400
We have three forces acting on a car: P, the weight and friction force.
For the motion with constant speed all forces need to be in equilibrium:
P-mg sin⁡α-μmg cos⁡α=0
P=mg(sin⁡α+μ cos⁡α)
P=1500(9.8)(sin⁡10+0.7 cos⁡10 )=13 kN.